# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v2

@@ Abstract
Large Language Models (LLMs) excel at providing information acquired during pretraining on large-scale corpora and following instructions through user prompts. This study investigates whether the quality of LLM responses varies depending on the demographic profile of users. Considering English as the global lingua franca, along with the diversity of its dialects among speakers of different native languages, we explore whether non-native English speakers receive lower-quality or even factually incorrect responses from LLMs more frequently. Our results show that performance discrepancies occur when LLMs are prompted by native versus non-native English speakers and persist when comparing native speakers from Western countries with others. Additionally, we find a strong anchoring effect when the model recognizes or is made aware of the user’s nativeness, which further degrades the response quality when interacting with non-native speakers. Our analysis is based on a newly collected dataset with over 12,000 unique annotations from 124 annotators, including information on their native language and English proficiency.
@@ 1 Introduction
English, as the global lingua franca, is predominant in large-scale text corpora used to train Large Language Models (LLMs) (Ziems et al., 2023; Zhang et al., 2023), including widely used datasets like CommonCrawl. These datasets are primarily tailored to an English-speaking audience located in the United States, and are mainly composed of privileged English dialects from wealthier educated urban zones (Talat et al., 2022; Ziems et al., 2023; Ryan et al., 2024; Gururangan et al., 2022). This biased training dataset composition permeates the LLM, resulting in models tailored to these English dialects (Santy et al., 2023; Hall et al., 2022). This highlights underlying design biases in LLMs, a phenomenon where design choices result in improved downstream performance for specific sub-populations (Santy et al., 2023). Consequently, their effectiveness considerably decreases when prompted in other languages or in underrepresented English dialects (Lai et al., 2023; Zhang et al., 2023; Bang et al., 2023; Ziems et al., 2023; Ryan et al., 2024).
Figure 1: An example prompt of a native and non-native English speaker and the corresponding output given by GPT4o. The desired output is C. The model selects the wrong answer for the non-native English speaker, while semantically the same message was conveyed.
@@
LLMs are highly sensitive to prompt formulations Beck et al. (2024); Chakraborty et al. (2023).  Ryan et al. (2024) show how models’ responses are tailored to Western English dialects, with prompt selection impacting LLMs’ preference tuning. Therefore, prompting models in other dialects can result in performance differences due to these design biases. Ziems et al. (2023) even provide a dataset covering multiple English dialects. However, unlike those studies focusing only on English dialects from English-speaking countries, our research also incorporates participants from countries where English is not an official language. We assess if word sensitivity in prompts disproportionately benefits native English speakers, leading to better model performance. In this case, the model has an inherent native language bias.
@@
In this paper, we find performance differences when LLMs are prompted by native versus non-native English speakers. More specifically, models often generate inaccurate responses for non-native speakers and rate the native prompts more positively than intended. We collect a dataset comprising over 12,000 unique prompts from native and non-native English speakers around the world, and demonstrate how different prompt formulations can lead to worse performance despite conveying the same message. An example of a prompt in our dataset is shown in Figure 1. Moreover, we find that these performance differences increase when comparing native English speakers from Western countries (US, UK, Canada) with other native and non-native English speakers. When the model recognizes or is informed about the user’s nativeness, a strong anchoring effect occurs, where model performance is substantially affected, leading to increased bias towards native English speakers.
@@
Our contributions are as follows: 1) We quantitatively and qualitatively analyze the performance of LLMs on objective and subjective classification tasks [^1] By subjective tasks, we mean classification tasks where the correct answer depends on the subjective interpretation as explained in Beck et al. (2024), as well as generative tasks, when prompted by native and non-native English speakers. 2) We investigate the impact on the LLM performance when the nativeness of the user is explicitly stated or inferred by the model. 3) We publish our multilingual instruction-tuning dataset and code used for the experiments [^2] https://github.com/manon-reusens/native_en_bias containing over 12,000 unique prompts from a diverse group of native and non-native English speakers worldwide, including translations of the prompts into eight different native languages.
@@ 2 Related work
Model Positionality and Design Bias. Model positionality, coined by Cambo and Gergle (2022), refers to the social and cultural position of a model, influenced by the stakeholders involved in its development, such as annotators and developers. This positionality affects the inclusivity of LLMs, as they evolve with certain biases that may disadvantage specific populations (Cambo and Gergle, 2022; Santy et al., 2023). Design biases arise when researchers make choices that improve model performance for specific sub-populations Santy et al. (2023). A notable example is the overrepresentation of English pretraining corpora, which leads to disproportionate performance improvements in English compared to other languages (Qin et al., 2023; Blasi et al., 2022; Joshi et al., 2020).
Effect of demographic background on LLM performance. Recent literature suggests that LLM performance on subjective tasks is influenced by the demographic attributes of the user (Beck et al., 2024; Santy et al., 2023). Moreover, when assigned a persona, LLMs reveal deep inherent stereotypes against various socio-demographic groups (Cheng et al., 2023; Gupta et al., 2023; Deshpande et al., 2023). For example, Gupta et al. (2023) show how ChatGPT3.5, when asked to solve a math question while adopting the identity of a physically disabled person, generates that it cannot answer the question, as a physically disabled person. Furthermore, Barikeri et al. (2021) demonstrate that LLMs can infer demographic attributes from dialog interactions. Additionally, research shows biases in favor of Western populations (Santy et al., 2023; Durmus et al., 2023). In model alignment literature, Ryan et al. (2024) show this similar bias within preference models and Gururangan et al. (2022) illustrate that even within a Western country like the US, GPT3 prefers the more privileged dialects. Finally, Ziems et al. (2023) have provided a cross-dialectal English dataset for countries with English as an official language. Building on these findings, we extend the research to include non-native English speakers, who use English dialects influenced by their native languages. Furthermore, while Gupta et al. (2023) assign a persona to the model, we analyze performance differences of LLMs both with and without explicitly informing the model about the user’s native language and thus with and without assigning a persona to the prompt writer.
@@ 3 Methodology
Given the sensitivity of LLMs to prompt formulation (Beck et al., 2024; Chakraborty et al., 2023), the diversity of English dialects (Ziems et al., 2023; Ryan et al., 2024), and alignment of models towards Western native English speakers (Ryan et al., 2024; Santy et al., 2023; Gururangan et al., 2022), we hypothesize that these design choices affect LLM performance when interacting with native versus non-native English speakers. Furthermore, we hypothesize that the specific English dialect learned may influence the responses from LLMs, particularly if the LLMs are optimized for English variants from the US, UK, and other predominantly English-speaking countries, as shown by Santy et al. (2023). We also anticipate that performance differences increase when a model is explicitly informed about the user’s nativeness. Given the LLM’s ability to infer demographic annotator information from text (Staab et al., 2023), we also expect the LLM to infer a user’s nativeness and that this correct inference leads to greater performance differences.
To test these hypotheses, we collected a new dataset containing both classification and generation tasks, along with information about the native languages of the annotators, as this is lacking in existing literature. An overview of our methodology and experimental setup is shown in Figure 2.
Figure 2: Methodology and experimental setup. The upper part of the figure shows the data collection steps. After gathering the different datasets, study participants annotated the examples. Then we validated them and used them as input to generate LLM responses. The lower part of the figure shows the evaluation phase. Before gathering the results, we got a new ground truth value for the dataset based on the original example for the generation tasks.
@@ 3.1 Dataset
We selected ten diverse datasets from various natural language instruction tasks [^3] https://github.com/allenai/natural-instructions (Mishra et al., 2022; Wang et al., 2022), covering classification (subjective and objective) and generation tasks. These tasks, representing typical LLM interactions, follow a standard instruction pattern and should not inherently favor native speakers. The tasks include paraphrasing, article generation based on a summary or title, sentiment analysis, natural language understanding, and multiple-choice answering.
@@ 3.2 Annotations
Annotators had to have a minimum English proficiency level equivalent to a high school or university degree. Each annotator worked on 20 to 240 examples. We gathered them through direct recruitment, opting for an open annotation process rather than an existing annotation platform to ensure high-quality annotations. All annotators were reimbursed at a rate of at least 12.11 euros per hour.
In addition to gathering self-reported linguistic information, i.e. the respective native language, level of English, and frequency of English use, we also collected information from native English speakers about how they learned English. To investigate our hypotheses, we categorize English speakers based on their nativeness and based on their learning contexts. The term Western native refers to native English speakers who learned English from native speakers from countries like the UK, US, Australia, or Canada. The term not Western native refers to all other annotators in our dataset.
@@ 4.1 Gathering LLM responses
Using gathered annotations, we conducted experiments with the chat-versions of well-established LLMs, as these are used in daily life. An overview of the checkpoints per model is shown in Appendix F. We included GPT3.5555https://openai.com/index/gpt-3-5-turbo-fine-tuning-and-api-updates/, GPT4o [^6] https://openai.com/index/hello-gpt-4o/, Haiku (Anthropic, 2024), Sonnet (Anthropic, 2024), using the appropriate APIs, and Qwen1.5 7B [^7] We ran the experiments for Qwen using A100 GPUs. (Bai et al., 2023) in line with the provided licenses and all consistent with the intended use. This set includes models of varying sizes, different performances, and from different developers, ensuring a diverse representation. Moreover, Qwen, developed by Chinese researchers, provides an interesting comparison in terms of design bias.
We experimented with various prompt schemata, structuring our methodological setup as follows. Firstly, we analyzed the results on our dataset without any modifications, comparing responses for native and non-native English speakers. Additionally, we distinguished between Western native English speaking and not Western native English speaking. Secondly, we hypothesize that the bias becomes more pronounced when adding explicit information on whether the annotator is a native or non-native English speaker, thereby analyzing whether an anchoring effect occurs. Anchoring is a term used for human cognitive bias indicating that a person might insufficiently change its estimates away from an initially provided value (Jones and Steinhardt, 2022; Tversky and Kahneman, 1974). This effect is demonstrated in LLMs by Jones and Steinhardt (2022), who found that code generation models modify their outputs to align with related solutions included in the prompt. Finally, we analyze the performance difference when letting the model first guess the nativeness of the annotator.
@@ 4.2 Evaluation
To measure the bias within the models, we look into the performance difference between two groups, i.e. native versus non-native and Western native versus not Western native. We measure these performance differences across classification tasks and generative tasks. Concretely, native bias measured for the classification tasks is defined as follows:
$\begin{split}\Delta_{\text{native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right),\psi\right)\\ \Delta_{\text{non-native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right),\psi\right)\end{split}$
@@
$\Delta_{native}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right),\mathcal{M}\left(\mathcal{T}\mid x_{\text{original}}\right)\right)$
$\begin{split}\Delta_{\text{non-native}}=&\phi(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right),\\ &\mathcal{M}\left(\mathcal{T}\mid x_{\text{original}}\right))\end{split}$
with native bias generative = $\Delta_{\text{native}}-\Delta_{\text{non-native}}$, template $\mathcal{T}$, user prompt $x$, model $\mathcal{M}$, performance metric $\phi$, and generated ground truth from the original prompt $\mathcal{M}\left(\mathcal{T}\mid x_{\text{original}}\right)$. The Western native bias can be similarly inferred.
Classification tasks. When assessing classification tasks, we focus on the accuracy of the predictions. We only consider classifications as correct if they follow the instructions correctly or if the correct classification can be determined automatically.
Generative tasks. In assessing the generative tasks, we include the following metrics: BLEU, ROUGE, BERT-score, and BART-score. These metrics are explained in detail in Appendix G.
The performance metrics for generative tasks are calculated using the LLM-generated gold answer as reference. This gold answer was generated using the original prompt in the dataset as displayed in Figure 2. We used this instead of the original gold answer in the dataset, given that optimal answers are model and context-dependent. By employing the LLM-generated gold answers, we align responses for each model.

### v3

@@ Abstract
Large Language Models (LLMs) excel at providing information acquired during pretraining on large-scale corpora and following instructions through user prompts. However, recent studies suggest that LLMs exhibit biases favoring Western native English speakers over non-Western native speakers. Given English’s role as a global lingua franca and the diversity of its dialects, we extend this analysis to examine whether non-native English speakers also receive lower-quality or factually incorrect responses more frequently. We compare three groups—Western native, non-Western native, and non-native English speakers—across classification and generation tasks. Our results show that performance discrepancies occur when LLMs are prompted by the different groups for the classification tasks. Generative tasks, in contrast, are largely robust to nativeness bias, likely due to their longer context length and optimization for open-ended responses. Additionally, we find a strong anchoring effect when the model is made aware of the user’s nativeness for objective classification tasks, regardless of the correctness of this information. This anchoring effect is a form of cognitive bias shown to be present in LLMs where the model is highly influenced by additional information. Our analysis is based on a newly collected dataset with over 12,000 unique annotations from 124 annotators, including information on their native language and English proficiency.
@@ 1 Introduction
English, as the global lingua franca, is predominant in large-scale text corpora used to train Large Language Models (LLMs) (ziems-etal-2023-multi; zhang-etal-2023-dont), including widely used datasets like CommonCrawl. These datasets are primarily tailored to an English-speaking audience located in the United States, and are mainly composed of privileged English dialects from wealthier educated urban zones (talat-etal-2022-reap; ziems-etal-2023-multi; ryan2024unintended; gururangan-etal-2022-whose). This biased training dataset composition permeates the LLM, resulting in models tailored to these English dialects (santy-etal-2023-nlpositionality; hall2022systematic). This highlights underlying design biases in LLMs, a phenomenon where design choices result in improved downstream performance for specific sub-populations (santy-etal-2023-nlpositionality). Consequently, their effectiveness considerably decreases when prompted in other languages or in underrepresented English dialects (lai-etal-2023-chatgpt; zhang-etal-2023-dont; bang-etal-2023-multitask; ziems-etal-2023-multi; ryan2024unintended).
(a) The desired output is C. This is an example prompt of an objective classification task.
(b) The desired output is 4. This is an example prompt of an subjective classification task.
Figure 1: Two example prompts of a native and non-native English speaker and the corresponding output given by GPT4o, where Annotator Prompt represents the placeholder for the annotations. For the objective task, the model selects the wrong answer for the non-native English speaker, while semantically the same message was conveyed. While Sentence B from the non-native speaker ("People can’t be eaten.") may seem different from Sentence A from the native speaker, it is a direct translation from the non-native speaker’s first language and conveys the same meaning from the non-native prompt writer’s perspective. This demonstrates how slight variations in phrasing, common among non-native speakers, can lead to misinterpretations or different model responses, despite semantic equivalence. For the subjective task, we see how the model estimates the native answer to be more positive than actually intended.
@@
LLMs are highly sensitive to prompt formulations beck-etal-2024-sensitivity; chakraborty-etal-2023-zero.  ryan2024unintended show how models’ responses are tailored to Western English dialects, with prompt selection impacting LLMs’ preference tuning. Therefore, prompting models in other dialects can result in performance differences due to these design biases. ziems-etal-2023-multi even provide a dataset covering multiple English dialects. However, unlike those studies focusing only on English dialects from native English-speaking countries, our research also incorporates participants from countries where English is not an official language. We assess if word sensitivity in prompts disproportionately benefits native English speakers, leading to better model performance. In this case, the model has an inherent native language bias.
@@
In this paper, we examine performance differences when LLMs are prompted by speakers from three groups: Western native (WN), non-Western native (NWN), and non-native (NN) English speakers. We find performance differences when LLMs are prompted by both NWN and NN versus WN speakers. More specifically, some models generate inaccurate responses for non-native speakers and rate the WN prompts more positively than intended. We also highlight how LLMs are more robust against this native bias on generative tasks. Moreover, we uncover deeply embedded bias within models towards native speakers for the classification tasks, as explicitly stating that a prompt writer is non-native leads to lower model performance compared to stating that the writer is native regardless of the correctness of this information. We collect a dataset comprising over 12,000 unique prompts from native and non-native English speakers worldwide and demonstrate how different prompt formulations can lead to worse performance despite conveying the same message. An example prompt from our dataset is shown in Figure 1.
@@
Our contributions are as follows: 1) Native bias analysis: We quantitatively and qualitatively analyze how LLM performance differs between native — both Western and non-Western— and non-native English speakers on objective and subjective classification tasks [^1] By subjective tasks, we mean classification tasks where the correct answer depends on the subjective interpretation as explained in beck-etal-2024-sensitivity, as well as generative tasks. 2) Novel Dataset: We publish our multilingual instruction-tuning dataset and code used for the experiments [^2] https://anonymous.4open.science/r/native_en_bias-EDC5/README.md containing over 12,000 unique prompts from diverse native and non-native English speakers, with translations into eight languages. 3) Innovative Data Collection: Our large-scale, structured annotation process across various tasks provides a comprehensive view of LLM responses from diverse user groups. 4) Novel Experimental Set-up: We propose a novel design evaluating the impact of informing the model about user nativeness, exploring whether it mitigates bias—an aspect not systematically studied before.
@@ 2 Related work
Model Positionality and Design Bias. Model positionality, coined by cambo2022model, refers to the social and cultural position of a model, influenced by the stakeholders involved in its development, such as annotators and developers. This positionality affects the inclusivity of LLMs, as they evolve with certain biases that may disadvantage specific populations (cambo2022model; santy-etal-2023-nlpositionality). Design biases arise when researchers make choices that improve model performance for specific sub-populations santy-etal-2023-nlpositionality. A notable example is the overrepresentation of English pretraining corpora, which leads to disproportionate performance improvements in English compared to other languages (qin-etal-2023-chatgpt; blasi-etal-2022-systematic; joshi-etal-2020-state).
Effect of demographic background on LLM performance. Recent literature suggests that LLM performance on subjective tasks is influenced by the demographic attributes of the user (beck-etal-2024-sensitivity; santy-etal-2023-nlpositionality). Moreover, when assigned a persona, LLMs reveal deep inherent stereotypes against various socio-demographic groups (cheng-etal-2023-marked; gupta2023bias; deshpande-etal-2023-toxicity). For example, gupta2023bias show how ChatGPT3.5, when asked to solve a math question while adopting the identity of a physically disabled person, generates that it cannot answer the question, as a physically disabled person. Furthermore, barikeri-etal-2021-redditbias demonstrate that LLMs can infer demographic attributes from dialog interactions. Additionally, research shows biases in favor of Western populations (santy-etal-2023-nlpositionality; durmus2023towards). In model alignment literature, ryan2024unintended show this similar bias within preference models and gururangan-etal-2022-whose illustrate that even within a Western country like the US, GPT3 prefers the more privileged dialects. Furthermore, hofmann2024ai illustrate how models show covert biases towards African American English speakers. Additionally, kantharuban2024stereotype show how LLMs express racially stereotypical recommendations regardless of whether the user explicitly or implicitly revealed their identity. Finally, ziems-etal-2023-multi have provided a cross-dialectal English dataset for countries with English as an official language. Building on these findings, we extend the research to include non-native English speakers, who use English dialects influenced by their native languages. Furthermore, while gupta2023bias assign a persona to the model, we analyze performance differences of LLMs both with and without explicitly informing the model about the user’s native language and thus with and without assigning a persona to the prompt writer. However, note that models providing different answers based on demographic background is not always problematic as noted in jin-etal-2024-implicit.
@@ 3 Methodology
Given the sensitivity of LLMs to prompt formulation (beck-etal-2024-sensitivity; chakraborty-etal-2023-zero), the diversity of English dialects (ziems-etal-2023-multi; ryan2024unintended), and alignment of models towards Western native English speakers (ryan2024unintended; santy-etal-2023-nlpositionality; gururangan-etal-2022-whose), we investigate whether LLMs exhibit bias in favor of native English speakers over non-native speakers. More specifically, we aim to answer the following research questions:
  1. 1.
    Do LLMs perform differently when prompted by native vs. non-native English speakers? And is there a performance difference for different groups of native English speakers?
  2. 2.
    Are certain tasks more prone to performance disparities between native and non-native speakers?
  3. 3.
    Which tasks, if any, remain robust to these differences?
  4. 4.
    Are these trends consistent across models, or do they vary by architecture?
  5. 5.
    Does explicitly providing information about a speaker’s nativeness amplify performance gaps?
To answer these research questions, we collected a new dataset containing both classification and generation tasks, along with information about the native languages of the annotators, as this is lacking in existing literature. An overview of our methodology and experimental setup is shown in Figure 2.
Figure 2: Methodology and experimental setup. The left part shows the data collection steps. After gathering the different datasets, study participants annotated the examples. Then we validated them and used them as input to generate LLM responses. The right part of the figure shows the evaluation phase, where we gathered the respective scores depending on the task.
@@ 3.1 Dataset
Our dataset was constructed including samples from ten diverse task datasets from various natural language instruction tasks [^3] https://github.com/allenai/natural-instructions (naturalinstructions; supernaturalinstructions), covering classification (subjective and objective) and generation tasks. These tasks, representing typical LLM interactions, follow a standard instruction pattern and should not inherently favor native speakers. The tasks include paraphrasing, article generation based on a summary or title, sentiment analysis, natural language understanding, multiple-choice answering, and review writing. This last task is the subjective classification task in our experiments. The different tasks provide varying levels of freedom in the dataset annotation tasks. This approach was explicitly chosen to have a range of more and less standardized annotation tasks where the level of freedom in prompt annotations varies depending on the underlying task. This way, our approach provides a comprehensive analysis of model performance.
@@ 3.2 Annotations
We required all annotators to have a minimum English proficiency level equivalent to a high school or university-level proficiency to establish a baseline, ensuring that performance differences stem primarily from dialectal variation rather than overall language proficiency. Each annotator worked on 20 to 240 examples. We gathered them through direct recruitment, opting for an open annotation process rather than an existing annotation platform to ensure high-quality annotations. All annotators were reimbursed at a minimum rate of 12.11 euros per hour.
In addition to gathering self-reported linguistic data—such as native language, English proficiency, and frequency of English use—we also collected information from native English speakers about how they acquired the language. This allows us to compare three groups: the non-native speakers (NN), Western native speakers (WN), and non-Western native speakers (NWN). The term Western native here refers to native English speakers who learned English from native speakers from countries like the UK, US, Australia, or Canada.
@@
We thus enforced strict quality control through the data collection phase to mitigate annotator variability through manual validation, removal of low-quality responses, and filtering examples with over 50% rejected annotations. This ensures that performance differences reflect linguistic or model-driven effects.
@@ 4.1 Gathering LLM responses
Using gathered annotations, we conducted experiments with the chat-versions of well-established LLMs, as these are used in daily life. An overview of the checkpoints per model is shown in Appendix F. We included GPT3.5555https://openai.com/index/gpt-3-5-turbo-fine-tuning-and-api-updates/, GPT4o [^6] https://openai.com/index/hello-gpt-4o/, Haiku (anthropic2024claude), Sonnet (anthropic2024claude), using the appropriate APIs, and Qwen1.5 7B [^7] We ran the experiments for Qwen using A100 GPUs. (bai2023qwen) in line with the provided licenses and all consistent with the intended use. This set includes models of varying sizes, different performances, and from different developers, ensuring a diverse representation. Moreover, Qwen, developed by Chinese researchers, provides an interesting comparison in terms of design bias.
To answer our predefined research questions mentioned in Section 3, we first ran our experiments for all models without any additional information. Next, to answer the last research question, we provided information about the nativeness of the prompt writer to the LLM. To see whether the LLM entails an inherent bias against native speakers, we included both correct and incorrect information.
@@ 4.2 Evaluation
To measure the bias within the models, we look into the performance difference between the native and non-native speaking groups. These performance disparities could contribute to allocational harms and representational harms, as defined by blodgett-etal-2020-language. Allocational harms can arise if non-native prompts result in systematically lower-quality responses, potentially affecting users’ access to accurate information, career guidance, or educational support. Similarly, representational harms may arise if certain English varieties are implicitly treated as less legitimate, reinforcing linguistic hierarchies and marginalizing speakers of underrepresented dialects.
We measure these performance differences across classification tasks and generative tasks. Concretely, native bias measured for the classification tasks is defined as follows:
$\Delta_{\text{native}}$ $=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right),\psi\right)$
$\Delta_{\text{non-native}}$ $=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right),\psi\right)$
@@
$\Delta_{\text{native}}$ $=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right)\right)$
$\Delta_{\text{non-native}}$ $=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right)\right)$
with native bias generative = $\Delta_{\text{native}}-\Delta_{\text{non-native}}$, template $\mathcal{T}$, user prompt $x$, model $\mathcal{M}$, and performance metric $\phi$. The Western native bias can be similarly inferred by splitting the native group into a Western native and non-Western native group.
Classification tasks. When assessing classification tasks, both objective and subjective classification tasks, we focus on the accuracy of the predictions. We only consider classifications as correct if they follow the instructions correctly or if the correct classification can be determined automatically.
Generative tasks. In assessing the generative tasks, we include the following metrics: fluency, coherence, and relevance (Bavaresco2024JUDGE_BENCH). All metrics were evaluated using a Likert scale: fluency was rated on a 3-point scale. Coherence and relevance were scored on a 5-point scale. Fluency is defined as the quality of the generated text in terms of grammar, spelling, etc. Coherence assesses the collective quality of the sentences. Finally, relevance refers to the inclusion of important content in the generated text. These definitions are based on the ones used in Bavaresco2024JUDGE_BENCH. The prompt templates used are shown in Appendix G. All results were rescaled to a range of 0 to 1 to ensure clarity. We evaluated the performance of the generative tasks using an LLM-as-a-judge approach, specifically leveraging Llama-3.3-70B-Instruct to assess each prompt’s output according to the three generative metrics mentioned earlier. For transparency, we have also included the exact evaluation prompts in Appendix G. To ensure reliability of the LLM-generated responses, we manually annotated 100 examples and observed a correlation of 81.3% with the model’s evaluations. The Cohen’s kappa score is 0.5564. However, given that the generation results are evaluated on a three- and five-point Likert scale, the correlation score is the most informative metric.